# Interpreting Equivariant Representations

### v1

@@ Abstract
Latent representations are used extensively for downstream tasks, such as visualization, interpolation or feature extraction of deep learning models. Invariant and equivariant neural networks are powerful and well-established models for enforcing inductive biases. In this paper, we demonstrate that the inductive bias imposed on the by an equivariant model must also be taken into account when using latent representations. We show how not accounting for the inductive biases leads to decreased performance on downstream tasks, and vice versa, how accounting for inductive biases can be done effectively by using an invariant projection of the latent representations. We propose principles for how to choose such a projection, and show the impact of using these principles in two common examples: First, we study a permutation equivariant variational auto-encoder trained for molecule graph generation; here we show that invariant projections can be designed that incur no loss of information in the resulting invariant representation. Next, we study a rotation-equivariant representation used for image classification. Here, we illustrate how random invariant projections can be used to obtain an invariant representation with a high degree of retained information. In both cases, the analysis of invariant latent representations proves superior to their equivariant counterparts. Finally, we illustrate that the phenomena documented here for equivariant neural networks have counterparts in standard neural networks where invariance is encouraged via augmentation. Thus, while these ambiguities may be known by experienced developers of equivariant models, we make both the knowledge as well as effective tools to handle the ambiguities available to the broader community.
@@ 1 Introduction
Figure 1: Visualizing rotated MNIST images via their equivariant representation (left) hides structure that becomes apparent using an invariant representation of the same latent codes (right).
Latent representations are used extensively in the interpretation and design of deep learning models. The latent spaces of VAEs (Kingma et al., 2019) and masked autoencoders (He et al., 2022), learned from great amounts of unlabeled data in an unsupervised or self-supervised manner, are prominent examples of powerful latent representations. These representations are used for e.g. molecular-, chemical- or protein discovery (Detlefsen et al., 2022); image generation, synthesis (Goodfellow et al., 2020) and segmentation (Kirillov et al., 2023); interpretability using visualization tools like T-SNE (Van der Maaten & Hinton, 2008), UMAP (McInnes et al., 2018) or PCA on latent embeddings; or counterfactual explainability (Papernot & McDaniel, 2018). It is natural to ask for such autoencoders – we use VAEs as a running example – to be equivariant. However, analysing equivariant latent features is no trivial task, and may yield misguided conclusions if done incorrectly.
@@
We contribute
an empirical demonstration of how naïve interpretation of equivariant latent spaces can lead to incorrect conclusions about data and decreased performance of derived models. We mathematically explain how analysis of equivariant latent representations needs to take the group action into account. We provide explicit tools for doing so using invariant projections of latent spaces. We evaluate their effect via widely encountered group actions on two widely used model classes: 1) A permutation equivariant variational autoencoder (VAE) representing molecular graphs acted on by node permutations, where we obtain isometric invariant representations of the data, and 2) an equivariant representations of a rotation-invariant image classifier, where we showcase random invariant projections as a general and efficient tool for providing expressive invariant representations.
Finally, we show that the ambiguity of equivariant latent representation also extends to completely standard deep learning models, where invariance is encouraged via augmentation. An empirical example shows that even for such models, latent representations display behavior similar to that observed in equivariant latent spaces. Thus, while these ambiguities might be known and avoided by experienced developers of equivariant models, pointing them out and providing tools to handle them is of high relevance to all users that encode implicit biases via either equivariance or augmentation.
2 Latent representations of equivariant models are ambiguous
In this section, we aim to review the mathematical tools needed to discuss the equivariant representation illustrated in Fig. 1. As above, let $h$ be a model with a latent feature embedding $f$:
[…]
We consider $\mathcal{Z}$ to be a $G$-equivariant representation. That is, the feature embedding $f$ is $G$-equivariant and either: $k$ is $G$-equivariant (e.g. as in an autoencoder or a segmentation network) or $k$ is a $G$-invariant predictor (e.g. as in a classification or regression task).
Typically, $f$ and $k$ are themselves complex models consisting of many layers – in an autoencoder setup, $f$ would be the encoder and $k$ the decoder. This architecture prototypes a typical classification or regression network built with a geometric prior in mind, and has been suggested as a general blueprint for deep learning architectures (Bronstein et al., 2021). However, the interpretation and utilization of the equivariant latent representation $\mathcal{Z}$ is non-trivial, as any two points $x_{1},x_{2}\in\mathcal{X}$, which are equal up to some group element $g\in G$, i.e. $x_{1}=g\cdot x_{2}$, will not necessarily map to the same latent representation – they may in fact map to latent representations $z_{1}=f(x_{1})$ and $z_{2}=f(x_{2})=g\cdot f(x_{1})$ which are far apart in the latent space. If standard latent space analysis methods are applied out of the box on such equivariant representations, this can create problems with downstream analysis, which typically relies on similar data points having nearby latent representations.
The modelling-assumption that two elements $x_{1}$ and $x_{2}$ are equivalent, if there exists a group element $g$ transforming one into the other is encapsulated by the quotient space $\mathcal{X}/G$, which is the set consisting of all orbits $G\cdot x=[x]=\{g\cdot x|g\in G\}$ (Bredon, 1972). While quotient spaces have been used extensively in classical approaches to statistics and machine learning with geometric priors, they often come with non-Euclidean structure and singularities, severely inhibiting the availability of tools for statistics, optimization, and learning (Feragen & Nye, 2020; Mardia & Dryden, 1989; Kolaczyk et al., 2020; Severn et al., 2022; Calissano et al., 2023). Equivariant models implicitly encode the structure of quotient spaces, while, from the viewpoint of implementation and optimization, Euclidean tools are available with all the computational advantages they may offer.
@@ 2 Latent representations of equivariant models are ambiguous
[…]
Where $\pi$ denotes the canonical projection. The induced feature embedding $f^{\prime}$ can be defined by via representatives $x$ from each orbit. That is,
@@
for every orbit $[x]\in\mathcal{X}/G$. The quotient latent representation $\mathcal{Z}$ comes with a quotient metric defining distances between orbits:
@@
This perfectly illustrates the problem: the equivariant latent representation $\mathcal{Z}$ is analyzed using arbitrary equivariant representatives $g_{1}\cdot z_{1}$ and $g_{2}\cdot z_{2}$ of the orbits $[z_{1}]$ and $[z_{2}]$, the relative distances between the latent codes $g_{1}\cdot z_{1}$ and $g_{2}\cdot z_{2}$, which would typically be used for downstream analysis, are not well defined, as they depend on the group elements $g_{1}$ and $g_{2}$ – and may be very different from the relative quotient distance $d([z_{1}],[z_{2}])$ within $\mathcal{Z}/G$. Thus, it is key that we enforce the modelling assumption before conducting any down-stream analysis of the data.
@@ 3 Invariant analysis of equivariant representations
As any equivariant function composed with an invariant function will be invariant, we can obtain invariant representations of our latent features by passing them through an invariant function. That is, we aim to find an invariant map $s\colon\mathcal{Z}\to\mathcal{Z}_{s}$, and use this to extract invariant latent features. An obvious invariant projection is the quotient projection $\pi\colon\mathcal{Z}\to\mathcal{Z}/G$ – but as discussed above, the quotient $\mathcal{Z}/G$ comes with severe limitations for analysis. Instead, we seek an invariant feature representation $s$ where the invariant latent space $\mathcal{Z}_{s}$ is Euclidean. Note that any such $s$ can necessarily be written as a composition of $\pi$ with another mapping $s^{\prime}$:
Proposition 1.
@@
[…]
As the map $s$ is chosen post hoc, a main challenge is how to choose $s$ and $\mathcal{Z}_{s}$ to retain signal from the quotient space $\mathcal{Z}/G$ while simplifying the analysis. Choosing $s(z)=0$ is obviously invariant, but will destroy any signal in $\mathcal{Z}$.
Prop. 1 (proof in the appendix) highlights one of the problems of picking an arbitrary invariant map $s$, as we will in general only be guaranteed that $\mathcal{Z}_{s}$ is more coarse than $\mathcal{Z}/G$. We refer to elements of $\mathcal{Z}$ as equivariant representations and to elements of $\mathcal{Z}_{s}$ as invariant representations. In general, an invariant representation $s(z)\in\mathcal{Z}_{s}$ does not uniquely identify the orbits of an element $z\in\mathcal{Z}$. We argue that the quality of the invariant $s$ depends on whether it satisfies:
Bijectivity:
If $s$ induces a bijection $s^{\prime}$ an immediate consequence is that all orbits of elements in $\mathcal{Z}$ can be uniquely identified by the corresponding invariant representation.
Isometry:
Equivariant/invariant models carry an inductive bias ensuring that elements that are equal up to the group action are handled similarly. Adhering to this modelling assumption, latent features of such elements should be close. When $s$ induces an isometric function $s^{\prime}$, then this will hold exactly for the invariant representations, which is isometric to the quotient representation – i.e. elements in $\mathcal{Z}$ which are close up to the action of a group element will also be close in $\mathcal{Z}_{s}$.
In a special case visited below, where $\mathcal{Z}_{s}$ and $\mathcal{Z}/G$ are isometric and $\mathcal{Z}_{s}\subset\mathcal{Z}$, we will speak of $\mathcal{Z}_{s}$ as being an isometric cross section. In cases where such a cross section exists, we argue, that the equivariant representation of the latent features should be mapped onto the cross section prior to any subsequent analysis.
@@ 3.1 Retrieving an isometric cross section: Latent graph representation
The resulting map $s$ is clearly invariant, since for all $\sigma\in S_{n}$ we have that $\sigma z$ and $z$ will have the same form when sorted. Also, if we let $\mathcal{Z}_{s}=\{s(z)\mid z\in\mathcal{Z}\}$, then $s$ is surjective by definition. These observations, combined with Proposition 1, allow us to show the following results implying that $s$ does indeed induce an isometric cross section $s^{\prime}\colon\mathcal{Z}/S_{n}\to\mathcal{Z}_{s}$:
Proposition 2.
Let $s\colon\mathcal{Z}\to\mathcal{Z}_{s}$ be the sorting function described above. Furthermore equip $\mathcal{Z}$ and $\mathcal{Z}_{s}$ with the Euclidean metric, and $\mathcal{Z}/S_{n}$ with the quotient metric. Then the induced $s^{\prime}\colon\mathcal{Z}/S_{n}\to\mathcal{Z}_{s}$ defined as in Proposition 1 is a bijection as well as an isometry. Finally $\mathcal{Z}_{s}$ is a convex cone.
The realization that $\mathcal{Z}_{s}$ is a convex is important, because it makes linear interpolation between elements of $\mathcal{Z}_{s}$ meaningful: Any point on the the line will be contained in the $\mathcal{Z}_{s}$ as well. A proof of Proposition 2 is included in the appendix.
3.2 When no isometric cross section exists: Random invariant linear projection

### v2

@@ Abstract
Latent representations are extensively used for tasks like visualization, interpolation, or feature extraction in deep learning models. This paper demonstrates the importance of considering the inductive bias imposed by an equivariant model when using latent representations as neglecting these biases can lead to decreased performance in downstream tasks. We propose principles for choosing invariant projections of latent representations and show their effectiveness in two examples: A permutation equivariant variational auto-encoder for molecular graph generation, where an invariant projection can be designed to maintain information without loss, and for a rotation-equivariant representation in image classification, where random invariant projections proves to retain a high degree of information. In both cases, the analysis of invariant latent representations proves superior to their equivariant counterparts. Finally, we illustrate that the phenomena documented here for equivariant neural networks have counterparts in standard neural networks where invariance is encouraged via augmentation.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Visualizing rotated MNIST images via their equivariant representation (left) hides structure that is apparent in an invariant representation of the same latent codes (right).
Latent representations are used extensively in the interpretation and design of deep learning models. The latent spaces of VAEs (Kingma et al., 2019) and masked autoencoders (He et al., 2022), learned from large amounts of unlabeled data unsupervised or self-supervised, are prominent examples of powerful latent representations. These representations are used for e.g. molecular-, chemical- or protein discovery (Detlefsen et al., 2022); image generation, synthesis (Goodfellow et al., 2020) and segmentation (Kirillov et al., 2023); semantic interpolation (Berthelot et al., 2018); interpretability using visualization tools such as T-SNE (Van der Maaten & Hinton, 2008), UMAP (McInnes et al., 2018) or PCA on latent embeddings; or counterfactual explainability (Papernot & McDaniel, 2018).
However, the analysis of such latent representations is no trivial task, and may yield misguided conclusions if done naively. Usually, deep neural networks are designed with specific geometric inductive biases in mind (Bronstein et al., 2021); an inductive bias can be imposed on a deep learning model by requiring the model to be invariant/equivariant to certain transformations of the data (Cohen & Welling, 2016; Puny et al., 2022) represented by a group of symmetries. A consequence of incorporating such a property in the modelling procedure is to make sure that datapoints which are similar are be mapped to similar outputs. This observation has consequences for any representation we might want to learn, as we will in general aim for any learned representation to be invariant to the transformation in question, while maximizing expressive power (Wang et al., 2024). These invariant representations can be obtained in many ways using for instance data augmentation or an inherently invariant model (Kwon et al., 2023) or by quotienting out nuisance parameters (Williams et al., 2021), however, in the case of autoencoders – in this paper we use VAEs as a running example – it is natural design these to be equivariant, and thus the question of interest becomes: How to we obtain an invariant representation using an equivariant model?
We contribute:
An empirical demonstration of how a naive interpretation of equivariant latent spaces can lead to incorrect conclusions about data and decreased performance of derived models. We mathematically explain how analysis of equivariant latent representations needs to take the group action into account. We provide explicit tools for doing so using invariant projections of latent spaces. We evaluate the effect of the suggested tools via widely encountered group actions on two widely used model classes: 1) A permutation equivariant variational autoencoder (VAE) representing molecular graphs acted on by node permutations, where we obtain isometric invariant representations of the data, and 2) an equivariant representations of a rotation-invariant image classifier, where we showcase random invariant projections as a general and efficient tool for providing expressive invariant representations. Finally, we show that the ambiguity of equivariant latent representation also extends to standard deep learning models, where invariance is encouraged via augmentation. An empirical example shows that even for such models, latent representations display behavior similar to that observed in equivariant latent spaces. Thus, while these ambiguities might be known and avoided by experienced developers of equivariant models, pointing them out and providing tools to manage them is important to all users that encode implicit biases via equivariance or augmentation.
Invariance and equivariance:
@@
2 Equivariant Models Enable Implicit Quotient Learning
The modelling-assumption that two elements $x_{1}$ and $x_{2}$ are equivalent, if there exists a group element $g$ transforming one into the other is encapsulated by the quotient space $\mathcal{X}/G$, which is the set consisting of all orbits $G\cdot x=[x]=\{g\cdot x|g\in G\}$ (Bredon, 1972). While quotient spaces have been used extensively in classical approaches to statistics and machine learning with geometric priors, they often come with non-Euclidean structure and singularities, severely inhibiting the availability of tools for statistics, optimization, and learning (Feragen & Nye, 2020; Mardia & Dryden, 1989; Kolaczyk et al., 2020; Severn et al., 2022; Calissano et al., 2024). Equivariant models implicitly encode the structure of quotient spaces, while, from the viewpoint of implementation and optimization, Euclidean tools are available with all the computational advantages they may offer.
@@ 2 Equivariant Models Enable Implicit Quotient Learning
${\mathcal{X}}$${\mathcal{Z}}$${\mathcal{X}/G_{\mathcal{X}}}$${\mathcal{Z}/G_{\mathcal{Z}}}$$\scriptstyle{f}$$\scriptstyle{\pi}$$\scriptstyle{\pi}$$\scriptstyle{f^{\prime}}$
Where $\pi$ denotes the canonical projection. The induced feature embedding $f^{\prime}$ can be defined via representatives $x$ from each orbit. That is,
@@
for every orbit $[x]\in\mathcal{X}/G$. The quotient space $\mathcal{Z}/G$ comes with a quotient metric defining distances between orbits:
@@
As the quotient spaces $\mathcal{X}/G$ and $\mathcal{Z}/G$ can exhibit strongly non-Euclidean geometry (Calissano et al., 2024), it is usually not feasible to fit a model mapping between these two spaces directly. However, by defining $f$ to be equivariant we do effectively learn a map $f^{\prime}$ between the quotient spaces, without having to deal with the issues that the non-Euclidean geometry of the quotient space causes.
3 Equivariant Latent Representations are Ambiguous
We have previously stated that, in general, we prefer invariant representations to equivariant representations, but why is that? This is since the interpretation and utilization of an equivariant latent representation is non-trivial, as any two points $x_{1},x_{2}\in\mathcal{X}$, which are equal up to some group element $g\in G$, i.e. $x_{1}=g\cdot x_{2}$, will almost certainly not map to the same latent representation – they may in fact map to latent representations $z_{1}=f(x_{1})$ and $z_{2}=f(x_{2})=g\cdot f(x_{1})$ which are far apart in the latent space. If standard latent-space analysis methods are applied out of the box on such equivariant representations, it can create problems with downstream analysis, which typically relies on similar data points having nearby latent representations. An obvious example of this would be clustering based on the latent representations which often would rely on the pairwise Euclidean distances between the latent codes in the creation of the clusters.
In Figure 2 we illustrate the problem on a simple example: the equivariant latent representation $\mathcal{Z}$ is analyzed using arbitrary equivariant representatives $g_{1}\cdot z_{1}$ and $g_{2}\cdot z_{2}$ of the orbits $[z_{1}]$ and $[z_{2}]$, the relative distances between the latent codes $g_{1}\cdot z_{1}$ and $g_{2}\cdot z_{2}$, which would typically be used for downstream analysis, are not well defined, as they depend on the group elements $g_{1}$ and $g_{2}$ – and may be very different from the relative quotient distance $d([z_{1}],[z_{2}])$ within $\mathcal{Z}/G$. Thus, it is key that we extend our modelling assumption, about which data symmetries are present, to our representation before conducting any down-stream analysis of the data. In practice this is done by considering an invariant representation.
Figure 2: An illustration of the effect of applying sorting to $\mathbb{R}^{2}$. In this case, we see, that $\mathbb{R}^{2}$ is mapped to $\mathcal{Z}_{s}=\{(x,y)\in\mathbb{R}^{2}|x\leq y\}$ (i.e. the blue area). We see, that in the $S_{2}$ equivariant representation depicted on the left-hand side the similar objects (e.g. the pentagons) are not guaranteed to be close. After having applied an invariant map (sorting), we see, that this shortcoming is taken care of.
@@ 4 Invariant Analysis of Equivariant Representations
As any equivariant function composed with an invariant function will be invariant, we can obtain invariant representations of our latent features by passing them through an invariant function. That is, our objective is to find an invariant map $s\colon\mathcal{Z}\to\mathcal{Z}_{s}$, and use this to extract invariant latent features. An obvious invariant projection is the quotient projection $\pi\colon\mathcal{Z}\to\mathcal{Z}/G$, but, as discussed above, the quotient $\mathcal{Z}/G$ comes with severe limitations for analysis. Instead, we seek an invariant feature representation $s$ where the invariant latent space $\mathcal{Z}_{s}$ is Euclidean. Note that any such $s$ can necessarily be written as a composition of $\pi$ with another mapping $s^{\prime}$ as illustrated by the following proposition:
Proposition 4.1.
@@
${\mathcal{Z}}$${\mathcal{Z}_{s}}$${\mathcal{Z}/G}$$\scriptstyle{s}$$\scriptstyle{\pi}$$\scriptstyle{s^{\prime}}$
We will refer to elements of $\mathcal{Z}$ as equivariant representations and to elements of $\mathcal{Z}_{s}$ as invariant representations.
A proof of this proposition can be found in Appendix A.1. As the map $s$ is chosen post hoc, a main challenge is how to choose $s$ and $\mathcal{Z}_{s}$ to retain the signal from the quotient space $\mathcal{Z}/G$ while simplifying the analysis. Choosing $s(z)=0$ is obviously invariant, but will destroy any signal in $\mathcal{Z}$. Prop. 4.1 highlights one of the problems of picking an arbitrary invariant map $s$, as we will in general only be guaranteed that $\mathcal{Z}_{s}$ is more coarse than $\mathcal{Z}/G$ and that the neighbourhood structure in $\mathcal{Z}_{s}$ may be substantially different from the quotient $\mathcal{Z}/G$. Ideally, we would pick a map $s:\mathcal{Z}\to\mathcal{Z}_{s}$, which in turn induces an isometric embedding $s^{\prime}:\mathcal{Z}/G\to\mathcal{Z}_{s}$. However, one might consider what to do in cases, where such an embedding does not exist. In these cases we might relax the requirement that distances between points should be preserved and only require that $s^{\prime}$ is a homeomorphism. If we can ensure that this is the case, then we will effectively have preserved the neighborhood structure of the quotient space, which is in itself a very powerful guarantee. Luckily, classical point set topology, e.g. (Munkres, 2000), provides us with the recipe for ensuring that this is true. For the sake of clarity we first define:
Definition 4.2.
Let $X$ and $Y$ be metric spaces with metrics $d_{X}$ and $d_{Y}$. Assume that the map $f:X\to Y$ is a homeomorphism. Then $f$ is an embedding. If $f$ has the property that:
$d_{X}(x_{1},x_{2})=d_{Y}(f(x_{1}),f(x_{2}))\text{ for all }x_{1},x_{2}\in X$ (2)
Then $f$ is an isometric embedding.
Definition 4.3.
Let $\mathcal{Z}$ and $\mathcal{Z}_{s}$ be topological spaces, and let $f:\mathcal{Z}\to\mathcal{Z}_{s}$ be a surjective map. Then $f$ is said to be a quotient map provided $U\subseteq\mathcal{Z}_{s}$ is open if and only if $f^{-1}(U)\subseteq\mathcal{Z}$ is open.
Proposition 4.4.
Let $s:\mathcal{Z}\to\mathcal{Z}_{s}$ be defined as in proposition 4.1. Then the induced map $s^{\prime}$ is continuous if and only if $s$ is continuous; $s^{\prime}$ is a quotient map if and only if $s$ is a quotient map; $s^{\prime}$ is a homeomorphism if and only if $s^{\prime}$ is a bijective quotient map.
A proof of Proposition 4.4 can be found in Appendix A.2. We can now use Proposition 4.4 as a guideline for how to select an invariant map $s$ that preserves the topology of the quotient space in our invariant representation. An illustration of the possible choices can be seen in Fig. 3, where an invariant isometry would yield an invariant latent representation most faithful to the original. Note that the hierarchy above illustrates what to take into account when picking an invariant map which respects the topology of the quotient space to as large a degree as possible, however, if there exists no embedding (i.e. homeomorphism) of the quotient space, then one might choose to sacrifice neighbourhood information in favour of picking a bijective map, as in general continuous- and quotient maps $s(z)\in\mathcal{Z}_{s}$ does not uniquely identify the orbits of an element $z\in\mathcal{Z}$.
Figure 3: An illustration of possible properties of the induced map $s^{\prime}:\mathcal{Z}/G\to\mathcal{Z}_{s}$ to prioritize when choosing an invariant projection $s:\mathcal{Z}\to\mathcal{Z}_{s}$. Each category of mappings preserves a different amount of structure of the latent space.
@@ 4.1 Retrieving an Isometric Cross Section: Latent Graph Representation
In a special case visited below, where $\mathcal{Z}_{s}$ and $\mathcal{Z}/G$ are isometric and $\mathcal{Z}_{s}\subset\mathcal{Z}$, we will speak of $\mathcal{Z}_{s}$ as being an isometric cross section. In cases where such a cross section exists, we argue, that the equivariant representation of the latent features should be mapped onto the cross section prior to any subsequent analysis, as this will be analogous to considering an isometric embedding.
@@
The resulting map $s$ is clearly invariant, since $\sigma z$ and $z$ will have the same form when sorted for all $\sigma\in S_{n}$. Also, if we let $\mathcal{Z}_{s}=\{s(z)\mid z\in\mathcal{Z}\}$, then $s$ is surjective by definition. These observations, combined with Proposition 4.1, allow us to show the following results implying that $s$ does indeed induce an isometric cross section $s^{\prime}\colon\mathcal{Z}/S_{n}\to\mathcal{Z}_{s}$:
Proposition 4.5.
Let $s\colon\mathcal{Z}\to\mathcal{Z}_{s}$ be the sorting function described above. Furthermore, equip $\mathcal{Z}$ and $\mathcal{Z}_{s}$ with the Euclidean metric, and $\mathcal{Z}/S_{n}$ with the quotient metric. Then the induced $s^{\prime}\colon\mathcal{Z}/S_{n}\to\mathcal{Z}_{s}$ defined as in Proposition 4.1 is a bijection and an isometry. Finally $\mathcal{Z}_{s}$ is a convex cone.
A proof of Proposition A.3 is included in Appendix A. The realization that $\mathcal{Z}_{s}$ is a convex is important, because it makes linear interpolation between elements of $\mathcal{Z}_{s}$ meaningful: Any point on the the line will be contained in the $\mathcal{Z}_{s}$ as well.
4.2 A Choice of Continuous Map: Random Invariant Linear Projections